Find the rotational kinetic energy for the next system if it rotates at a rate of 10 revolutions per second.
4.00k4 y = 3.00m m2 2.0oka y = -2.00M m3 3.00 kg = -4.00m
Answer #1

2 KE = 1/ 1 w2 Is Mergh Masu - Mg 3² I= (4) (32) & (2) (2²) + (3) (42) I= 92 kg m² W= 10 x 215 rad Is = 62.83 rad 18 Ke = 18